W19- Chap Chapě Get He W19- W195 yPost Oka Knda wom 12 pts D Beth othand (A Busir “ HP-See whats Hot G ■Hp Games Simported From Mr. Espngs Math Sit Tub D Question 9 Match the acronym on the left with its correct formula/equation. Formulas may be used more than once or not at ait Sup that h availa NOPAT BIT(1-Tax Rate) FCF ·NOPAT-(Non-Cash le Strnt ▼| EBIAT [Choose I [ Choose 1 -NOPAT-INon-Cash Inc Stmt items) Chg in NWC)-CapEx (Chg in Regd Cash) EBIT(1-Tax Rate) Chg in RE-NI-DIV EBITDA1Tax Rate) Elrl p1ri+p2 12 p33.nr Re-Ri+B(Rm-Rf) Chg in NWO CAPM WdRd(1-Tax Rate)+WeRe . Previous -Chg in AR+ Chg in INV Chg in AP
Answer #1

1: NOPAT = EBIT*(1-Tax)

(NOPAT is net operating profit after tax. EBIT is operating profit. Hence NOPAT is computed as EBIT*(1-Tax))

2: FCF = NOPAT- (Non cash items)- Change in NWC- Capex- Change in cash

(Free cash flow refers to cash flow available after changes in working capital, fixed assets expenses and cash balance)

3: EBIAT= EBIT*(1-Tax)

(EBIAT is EBIT after tax)

4: Change in NWC = Chg in AR+ Chg in Inventory -Chg in AP

(This is difference between change in current assets and current liabilities)

5: WACC = WD*Rd*(1-Tax)+ We*Re

(WACC is the sum of weights*Cost of debt/equity)

6: CAPM : RE= Rf+ Beta*(Rm-Rf)

As per CAPM, Cost of equity is risk free rate+ Beta*Risk premium